OpenStudy (anonymous):

y^2+3-12/y^2+10y+24=

OpenStudy (anonymous):

What are you meant to do? Factor? Solve? Is this thing equal to something?

OpenStudy (anonymous):

factor and solve

OpenStudy (anonymous):

I think there might be something wrong with your equation. I see it as you asking me to factor the left-hand side and solve this:\[y^2+3-\frac{12}{y^2}+10y+24=0\] Or are you asking for,\[\frac{y^2+3y-12}{y^2+10y+24}\]Can you check please? People will read what you've written in your question and assume it's the first thing I wrote out. Also, I assumed in the second that you forgot to place a y after the 3.